# http://dimadamen.github.io/OSNOM

### v1

@@ 4.2 Experimental setup
Baselines. As no prior works have attempted the OSNOM task, we compare LMK against four baselines:
@@
  • EgoLoc [24]: we adapt this SOTA VQ3D approach to OSNOM, to handle multiple objects. We use the same masks, features, 3D scene and lifting for fair and direct comparison. EgoLoc’s weighted averaging over all past observations fails for OSNOM because objects change position, so instead we take the most recent match.
Implementation details. For appearance features $\Psi$ we use a DINO-v2 [29] pre-trained model. We crop each mask, scale to $224\times 224$ and pass to the backbone. We ablate the choice of features. We set $\alpha=10$, $\gamma=100$, $\beta_{L}=13$ and $\beta_{V}=2$ (chosen on the validation set). We compute meshes in advance, which takes 5 hours on average on one 2080Ti per video. Then for online tracking, DINOv2 operates at 30 FPS and lifting-to-3D at 200FPS on one P100. LMK runs at 1000fps on a single CPU core.
@@ 4.3 Results
Figure 3: OSNOM results. PCL of LMK compared to baselines.
Figure 3: OSNOM results. PCL of LMK compared to baselines.
@@
Results on the OSNOM task for LMK, compared to the baselines OSL, OSOM, Random Matching and EgoLoc, are shown in Figure 4. The average PCL (y-axis) over the whole dataset is reported for each 5s evaluation interval (shown on the x-axis), with standard deviation shaded. We show performance over both short (0-60 seconds) and long (1-12 minutes) timescales. Over time, the complexity of matching observations increases as more objects are being interacted with and tracked. This is reflected in performance decreasing for all methods over time.
LMK presents a significant improvement over all baselines. Compared to EgoLoc, it achieves a 39% average improvement tracking up to 1 minute, and 25% from 1 to 12 mins. This is because LMK tracks across consecutive frames, being more robust to variations in the appearance of objects due to changes in orientation or occlusion, and utilising 3D locations in the matching. The rapid drop in performance in OSOM and OSL shows the challenge of egocentric footage, where the constantly moving person causes objects to go out of view frequently. When objects are tracked until they are in-view (OSL baseline), performance goes to zero just after 20s, showing that objects are quickly lost from sight. The OSOM baseline shows that only considering objects within the camera’s field of view, without 3D world coordinates and object permanence, is insufficient for the OSNOM task (is worse than random).
@@
4.4 LMK Ablation
Accuracy at different radii. All our experiments set the PCL threshold, $R=30\mathrm{cm}$. Figure 9 also shows results when this is increased to $R=60\mathrm{cm}$ and $R=90\mathrm{cm}$, which are also visualised in 3D to show their relative difficulties. As expected, PCL increases as $R$ increases.
Weighting visual appearance and location. LMK uses the hyperparameters $\beta_{V}$ (Eq 4) and $\beta_{L}$ (Eq 5) for relative weighting of visual and location similarities when assigning new observations to tracks. We select these based on best validation set performance averaged over timescales. Figure 9(a) shows validation set performance when fixing the chosen $\beta_{V}=2$ and varying $\beta_{L}$. Figure 9(b) fixes $\beta_{L}=13$ and varies $\beta_{V}$. Both hyperparameters are relatively stable, most likely due to the scaling by appropriate distributions (Cauchy and Exponential).
Track visual appearance history. Figure 9(c) ablates $\gamma$ over the validation set – the number of recent features averaged for visual representation. Best results are obtained with $\gamma=100$, with worse results for small / large values of $\gamma$, with performance relatively stable even down to only one observation.
Figure 7: Evaluation thresholds. LMK when increasing the PCL threshold $R$ the maximum distance between predicted and ground truth 3D locationsconsidered successful. Visualisations show the regions covered by $R=30\mathrm{cm}/60\mathrm{cm}/90\mathrm{cm}$ volumes in blue, centered on the counter.
@@ 4.4 LMK Ablation
Figure 7: Evaluation thresholds. LMK when increasing the PCL threshold $R$ the maximum distance between predicted and ground truth 3D locationsconsidered successful. Visualisations show the regions covered by $R=30\mathrm{cm}/60\mathrm{cm}/90\mathrm{cm}$ volumes in blue, centered on the counter.
Figure 8: Visual feature choice of a DINO-v2, CLIP or ImageNet (ViT).
Figure 9: Detections. LMK on both visual and location features when using VISOR annotations vs using detections from [40].
@@
Object dimension. In our experiments, we use a fixed $R=30cm$. As objects differ in size, one might argue that matching $R$ to the object size is more reasonable. In Figure 13 we use an adapted $R$ that matches the object dimension per example. Results are extremely similar to the default $R=30cm$.
Effect of objects going out-of-view. We investigate the effect of a track going out- then back in-view (i.e. reappearing) within 10 minutes (Figure 13). LMK, matching using 3D locations, shows considerable performance improvement.
Figure 11: Object radius. LMK when using approximating objects as spheres in 3D and using object radius for PCL threshold R.
Figure 12: LMK for spatial cognition. Number of objects correctly located by LMK, separately by combinations of (In-reach, Out-of-reach) and (In-sight, Occluded, Out-of-view).
Figure 11: Object radius. LMK when using approximating objects as spheres in 3D and using object radius for PCL threshold R.
Figure 12: LMK for spatial cognition. Number of objects correctly located by LMK, separately by combinations of (In-reach, Out-of-reach) and (In-sight, Occluded, Out-of-view).
5 Conclusion
In this paper, we introduced the task of “Out of Sight, Not Out of Mind” (OSNOM) for egocentric video with partial object observations. It evaluates 3D tracking performance of active objects when they are both in- and out-of-sight. We introduced Lift, Match and Keep (LMK), a method which lifts partial 2D observations in camera coordinates to 3D world coordinates, matches them over time using visual appearance and 3D location, and keeps them in mind when they go out of sight. Results on long videos from EPIC-Kitchens show LMK delivers good results over both short ($64\%$ up to 1 minute) and long ($37\%$ for 1-12 minutes) timeframes, and that maintaining 3D world location is critical when objects go out-of-view. For future work, we will investigate whether LMK can help track objects that undergo state changes, and explore shared 3D object tracks between multiple ego- and exo-centric cameras.
Acknowledgments. Research at Bristol is supported by EPSRC Fellowship UMPIRE (EP/T004991/1) & PG Visual AI (EP/T028572/1). We particularly thank Jitendra Malik for early discussions and insights on this work. We also thank members of the BAIR community for helpful discussions. We acknowledge the use of University of Bristol’s Blue Crystal 4 (BC4) HPC facilities.
Appendix
Appendix 0.A Estimating error in the 3D projection
In Sec 4.1, we estimate the error in 3D locations, through comparing projections of static objects from multiple viewpoint. Figure 3 in the paper presented the findings – showcasing that the mean error is 3.5cm with 96% of all errors within 10cm. We here describe the data used to report this figure.
We randomly selected 207,277 pairs of frames from our dataset, covering correspondences between 10 static objects across 5 different kitchens/environments. These were manually selected as multiple frames with masks of the same object, at distinct times, and from different viewpoints. We avoid masks that are partially occluded by another object or by the camera’s field-of-view (i.e. not fully in view) as these projections are likely to differ due to the occlusion of part of the mask. As the chosen pairs of masks showcase the same static object, their 3D locations should perfectly match. Any differences in their 3D location can be used to measure the error in the 3D projection, which we use as ground truth locations.
Figure 14: LMK Results for Moved vs Stationary objects with respect to the environment.
Appendix 0.B Moved vs. Stationary objects
Section 3.3 also provides a definition of objects which have either moved or remained stationary with respect to the environment. Figure 14 shows PCL results using a movement threshold of $\epsilon=30\mathrm{cm}$. The performance of stationary objects is on average $35\%$ better than that of objects which have moved. Objects are more visually different after a move (e.g. different orientation or lighting).
Appendix 0.C Failure cases
@@
Appendix 0.D Future Directions

### v2

@@ 4.2 Experimental setup
Baselines. As no prior works have attempted the OSNOM task, we compare LMK against three naive baselines and three previous works adapted to the OSNOM task:
@@
  • ByteTrack [60]: a strong, recent 2D multi-object tracking method, widely used as a baseline [57, 23, 33]. Objects are tracked in 2D and then lifted in 3D using our lifting approach for evaluation.
  • EgoLoc [26]: we adapt this SOTA VQ3D approach to OSNOM, to handle multiple objects. We use the same masks and lifting for fair comparison. EgoLoc’s weighted averaging over all past observations fails for OSNOM because objects change position, so instead we take the most recent match.
  • IT3DEgo [61]: As this paper uses ground truth depth which is not available in our RGB sequences, we instead run the 2D tracking using their public code, then lift the tracked objects to 3D using our approach (Sec 3.1).
Figure 3: OSNOM results. PCL of LMK compared to baselines.
Implementation details. For appearance features $\Psi$, we use a DINO-v2 [31]. We crop each mask, scale to $224\times 224$ and pass to the backbone. We set $\alpha=10$, $\gamma=100$, $\beta_{L}=13$ and $\beta_{V}=2$ (chosen on the validation set). We compute meshes in advance, which takes 5 hours on average on one 2080Ti per video. Then for online tracking, DINOv2 operates at 30 FPS and lifting-to-3D at 200FPS on one P100. LMK runs at 1000fps on a single CPU core.
@@ 4.3 Results
Results on the OSNOM task for LMK, compared to the baselines, are shown in Figure 3. The average PCL (y-axis) over the whole dataset is reported for each 5s evaluation interval (shown on the x-axis), with standard deviation shaded. We show performance over both short (0-60 seconds) and long (1-12 minutes) timescales. Over time, the complexity of matching observations increases as more objects are being interacted with and tracked. This is reflected in performance decreasing for all methods over time.
LMK presents a significant improvement over all baselines. The rapid drop in performance in OSOM and OSL shows the challenge of egocentric footage, where the constantly moving person causes objects to go out of view frequently. When objects are tracked as long as in-view (OSL baseline), performance goes to zero just after 20s, showing that objects are quickly lost from sight. The OSOM baseline shows that only considering objects in-view, without 3D world coordinates and object permanence, is insufficient for the OSNOM task (is worse than random). LMK significantly outperforms ByteTrack, EgoLoc, and IT3DEgo. ByteTrack and IT3DEgo rely on 2D frames, while EgoLoc loses tracking quickly by comparing to initial appearances. In contrast, LMK tracks across consecutive frames, handling appearance changes from orientation or occlusion and leveraging 3D locations for robust matching.
4.4 LMK Ablation
@@
Figure 5: Evaluation thresholds. LMK when increasing the PCL threshold $R$ the maximum distance between predicted and ground truth 3D locations considered successful.
Figure 6: Detections. LMK and EgoLoc baseline [26] on both visual and location features (V+L) when using VISOR annotations vs using object detections from [43].
Figure 7: LMK for spatial cognition. Number of objects correctly located by LMK for each (In-reach, Out-of-reach) and (In-sight, Occluded, Out-of-view) combination.
@@
Accuracy at different radii. All our experiments set the PCL threshold, $R=30\mathrm{cm}$. Figure 5 also shows results when this is set to $R=10/20/60/90/120\mathrm{cm}$. As expected, PCL increases as $R$ increases.
We also investigate the ability of LMK to track objects going out- then back in-view (i.e. reappearing) within 10 minutes (Figure 8). LMK, matching using 3D locations, shows considerable performance improvement.
@@
We include examples of failure cases in Appendix D.
5 Conclusion
In this paper, we introduced the task of “Out of Sight, Not Out of Mind” (OSNOM) for egocentric video with partial object observations. It evaluates 3D tracking performance of active objects when they are both in- and out-of-sight. We introduced a very strong baseline: Lift, Match and Keep (LMK), a method which lifts partial 2D observations in camera coordinates to 3D world coordinates, matches them over time using visual appearance and 3D location, and keeps them in mind when they go out of sight. Results on long videos from EPIC-Kitchens show LMK delivers good results over both short ($64\%$ after tracking for 1 minute) and long ($37\%$ after 10 minutes) timeframes, and that maintaining 3D world location is critical when objects go out-of-view. LMK outperforms recent works, strong 2D tracks and naive baselines by a big margin. For future work, we will investigate whether LMK can help track objects that undergo state changes, and explore shared 3D object tracks between multiple ego- and exo-centric cameras.
Acknowledgments. Research at Bristol is supported by EPSRC Fellowship UMPIRE (EP/T004991/1) & PG Visual AI (EP/T028572/1).We particularly thank Jitendra Malik for early discussions and insights on this work. We also thank members of the BAIR community and the MaVi group at Bristol for helpful discussions. This project acknowledges the use of University of Bristol’s Blue Crystal 4 (BC4) HPC facilities.
Appendix A Video examples
Sample results are present on the project’s webpage: http://dimadamen.github.io/OSNOM. The video shows predicted object locations, over time, in 4 sampled clips from the evaluation EPIC-KITCHENS videos. We show the mesh of the environment, along with coloured neon dots representing the active objects that we lift and track in 3D. The videos also show the estimated camera position and direction throughout the video along with the corresponding egocentric footage.
In each case, the clip shows object locations predicted when they are in-sight, when they are out-of-view as well as when they are moving in-hand. Selected examples also show objects picked up / returned to fridge or cupboard highlighting the complexity of spatial cognition from egocentric videos.
Appendix B Estimating error in the 3D projection
In Section 4.1, we estimate the error in 3D locations, through comparing projections of static objects from multiple viewpoint. Figure 3 in the paper presented the findings – showcasing that the mean error is 3.5cm with 96% of all errors within 10cm. We here describe the data used to report this figure.
We randomly selected 207,277 pairs of frames from our dataset, covering correspondences between 10 static objects across 5 different kitchens/environments. These were manually selected to find multiple frames with masks of the same object, at distinct times, and from different viewpoints. We avoid masks that are partially occluded by another object or by the camera’s field-of-view (i.e. not fully in view) as these projections are likely to differ due to the occlusion of part of the mask. As the chosen pairs of masks showcase the same static object, their 3D locations should perfectly match. Any differences in their 3D location can be used to measure the error in the 3D projection, which we use as ground truth locations.
Figure 11: LMK Results for Moved vs Stationary objects with respect to the environment. We used a movement threshold of $\epsilon=30\mathrm{cm}$
@@ Appendix B Estimating error in the 3D projection
Figure 13: Object radius. LMK when approximating objects as spheres in 3D and using object radius for PCL threshold R.
@@
Appendix C Additional Ablations
Moved vs. Stationary objects. Section 3.3 also provides a definition of objects which have either moved significantly within the environment or remained relatively within a small section of the environment. We use a movement threshold of $\epsilon=30\mathrm{cm}$ to separate large from small motions. Figure 11 shows PCL results showing the objects that remain relatively stationary can be tracked on average $35\%$ better than that of objects which have moved significantly within the space. Objects are more visually different after a move (e.g. different orientation or lighting).
Object size. In our experiments, we use a fixed $R=30cm$. As objects differ in size, one might argue that matching $R$ to the object size is more reasonable. In Figure 13 we use an adapted $R$ that matches the object dimension per example. Results are very similar to the default $R=30cm$, showcasing that fixed versus dynamic $R$ do not change the tracking capabilities.
Weighting visual appearance and location. LMK uses the hyperparameters $\beta_{L}$ (Eq 4) and $\beta_{V}$ (Eq 5) for relative weighting of visual and location similarities when assigning new observations to tracks. We select these based on best validation set performance averaged over timescales. Figure 14(a) shows validation set performance when fixing the chosen $\beta_{V}=2$ and varying $\beta_{L}$. Figure 14(b) fixes $\beta_{L}=13$ and varies $\beta_{V}$. Both hyperparameters are relatively stable, most likely due to the scaling by appropriate distributions (Cauchy and Exponential).
Track visual representation. Figure 14(c) ablates $\gamma$ over the validation set – the number of recent features averaged for visual representation of a track. Best results are obtained with $\gamma=100$, with worse results for small / large values of $\gamma$, with performance relatively stable even down to only one observation.
Appendix D Failure cases
@@
Appendix E Future Directions